OpenStudy (anonymous):

Solve the recurrence relation \(a_n = 2^na_{n-1}\) with \(a_0 = 1\)

OpenStudy (anonymous):

I try \(a_n = 2^n a_{n-1} = 2^n(2^{n-1}a_{n-2}) = 2^{2n-1}a_{n-2} \) Is that a step in the right direction?

OpenStudy (michele_laino):

If we continue your equalities, we will get: \[a _{n}=2^{n}a _{n-1}=2^{n}2^{n-1}a _{n-2}=...2^{n}2^{n-1}2^{n-2}...2^{1}a _{0}\]

OpenStudy (michele_laino):

so \[a _{n}=2^{(1+n)\frac{ n }{ 2 }}\]

OpenStudy (michele_laino):

the formula above makes sense, because: if n=0, then a_0=1 if n=1, then a_1=2 if n=3, then a_3=2^6=64 and so on... @calyne

OpenStudy (anonymous):

amazing thank you

OpenStudy (michele_laino):

thank you!